# Characterizing rational homogeneous spaces via $\Csuperscript\C\C^{*}$-actions

### v1

@@ 4. Induced actions on families of rational curves
$(X,L)$ is a smooth polarized pair, endowed with an equalized ${\mathbb{C}}^{*}$-action, and admitting a beautiful family of rational curves ${\mathcal{M}}$, of $L$ degree $d$, with evaluation morphism $q:{\mathcal{U}}\to{\mathcal{M}}$ (see Section 2.7).
@@ 4.1. Action on ${\mathcal{M}}$
Let $C$ be a curve such that $[C]\in{\mathcal{M}}$; the point $[C]$ is fixed by the induced action if and only if $C$ is a ${\mathbb{C}}^{*}$-invariant curve, that is, either $C\subset Y$ with $Y\in{\mathcal{Y}}$ or $C$ is the closure of an orbit of the action. The latter curves will be called bridges of the family ${\mathcal{M}}$ (with respect to the ${\mathbb{C}}^{*}$-action). If a bridge $C$ has sink and source in fixed components $Y,Y^{\prime}$, we say that the bridge $C$ links $Y$ and $Y^{\prime}$. Note that in this case, by Lemma 2.3, $\mu_{L}(Y)\not=\mu_{L}(Y^{\prime})$, hence $Y\not=Y^{\prime}$.
@@
Given a non-fixed point $[C]\in{\mathcal{M}}$ we construct a rational ruled surface $S_{C}$ as follows: we take the closure of the orbit of $[C]$, $\overline{{\mathbb{C}}^{*}[C]}$, which is a rational curve, we take its normalization $\nu:{\mathbb{P}}^{1}\to\overline{{\mathbb{C}}^{*}[C]}$ and we consider the fiber square:
@@
By construction, the surface $S_{C}$ is of the form $\mathbb{F}_{e}:={\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(e))$ for some $e\geq 0$, and is endowed with a ${\mathbb{C}}^{*}$-action with two invariant fibers $F_{-}$, $F_{+}$, which are mapped via $\overline{q}$ to the limit curves of $C$.
@@
Note that $\overline{q}(S_{C})$ has dimension two, so $\overline{q}$ can contract at most one curve in $S_{C}$, otherwise we would have a one-dimensional family of curves of ${\mathcal{M}}$ passing by two points of $X$; this, by Bend and Break (see [7, Proposition 3.2]) would contradict the fact that ${\mathcal{M}}$ is an unsplit family. If the curve $C$ contains a fixed point of the action, then $S_{C}$ contains a section which is contracted by $\overline{q}$.
@@
We will now use Construction 1 to prove that the equalization of the action on $X$ implies the equalization of the the induced ${\mathbb{C}}^{*}$-action on ${\mathcal{M}}$.
@@
If $e=0$, we consider the ${\mathbb{C}}^{*}$-action on the space of global sections $\mathop{\rm H}\nolimits^{0}({\mathbb{P}}^{1},{\mathcal{E}})$, that has two linearly independent eigenvectors, providing two ${\mathbb{C}}^{*}$-invariant surjective morphisms ${\mathcal{E}}\to{\mathcal{O}}_{{\mathbb{P}}^{1}}$, that correspond to two ${\mathbb{C}}^{*}$-invariant sections of $S_{C}\to{\mathbb{P}}^{1}$, denoted by $C_{0},C_{1}$. We conclude noting that, by Remark 4.3, $\overline{q}$ cannot contract both $C_{0}$ and $C_{1}$.
@@
We finish the section describing the possibilities for the surface $S_{C}$ and the limit curves $C_{-},C_{+}$.
@@
    $e=1$, one among $C_{-}$ and $C_{+}$ is fixed, and the other is a bridge.
@@
Set ${\mathcal{E}}:={\mathcal{O}}_{{\mathbb{P}}^{1}}(1)\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(e+1)$ and let $a_{-},b_{-}$ (resp. $a_{+},b_{+}$) be the weights of the action at ${\mathcal{E}}_{[C_{-}]}$ (resp. at ${\mathcal{E}}_{[C_{+}]}$). By [26, Lemma 2.16], up to renaming and normalizing them, we have $a_{-}=0,a_{+}=1$ and $b_{+}=b_{-}+e+1$. Moreover, since the action on $F_{-}$ and $F_{+}$ is equalized we have $|b_{-}|\leq 1$ and $0\leq b_{+}\leq 2$. We also note that the fiber $F_{-}$ (resp. the fiber $F_{+}$) is a fixed point component if and only if $b_{-}=0$ (resp. $b_{+}=1$). Therefore we have the following possibilities, which lead to the cases listed in the statement.
@@ 00 $2$ $1$ fixed orbit
If $C$ contains a fixed point $y$, then every fiber of $S_{C}$ over ${\mathbb{P}}^{1}$ contains a point mapping to $y$; in other words, $S_{C}$ contains a section over ${\mathbb{P}}^{1}$ contracted via $\overline{q}$ to $y$. Since $C$ is not fixed, we may conclude that $e>0$. ∎
@@
Figure 1. The possible cases in Proposition 4.6
@@
  • In case 1), either $C_{+}$ links $Y$ to a component $Y^{\prime}$ of $L$-weight $\mu_{L}(Y)+d$, or $C_{-}$ links $Y$ to a component $Y^{\prime}$ of $L$-weight $\mu_{L}(Y)-d$.
@@ 4.2. Action on ${\mathcal{U}}$
Let $u\in{\mathcal{U}}$ be a non fixed point, and let $G_{u},G_{p(u)},G_{q(u)}\subset{\mathbb{C}}^{*}$ be the isotropy subgroups of $u\in{\mathcal{U}}$, $p(u)\in{\mathcal{M}}$, $q(u)\in X$. Clearly we have that $G_{u}\subset G_{p(u)}\cap G_{q(u)}$. Since the closure of ${\mathbb{C}}^{*}u$ cannot be contracted by both $p$ and $q$, we have that either $G_{p(u)}$ or $G_{q(u)}$ are finite. Moreover, since the actions of ${\mathbb{C}}^{*}$ on $X$ and ${\mathcal{M}}$ are equalized, we have that both $G_{p(u)}$ or $G_{q(u)}$ are trivial, and we conclude that $G_{u}$ is trivial as well. ∎
@@
Since $p$ is equivariant, the image of a fixed point in ${\mathcal{U}}_{Y}$ must be a point $[C]$ such that either $C\subset Y$ or $C$ is a bridge. In other words, the fixed point locus of ${\mathcal{U}}_{Y}$ can be written as a disjoint union of (a priori reducible) closed subsets
@@
Note that ${\mathcal{U}}_{Y}=\bigsqcup_{y\in Y}{\mathcal{M}}_{y}$, and that every ${\mathcal{M}}_{y}$ is ${\mathbb{C}}^{*}$-invariant. In particular, every orbit in ${\mathcal{U}}_{Y}$ is contained in ${\mathcal{M}}_{y}$ for some $y\in Y$. The following Lemma, which is a straightforward consequence of Corollary 4.7, describes the ${\mathbb{C}}^{*}$-orbits on ${\mathcal{U}}_{Y}$:
@@
If  ${\mathcal{U}}_{Y}^{i}$ (for $i=-,0,+$) contains an extremal fixed point component, then it is irreducible, and meets ${\mathcal{M}}_{y}$ for every $y$.
@@
Assume, without loss of generality, that ${\mathcal{U}}_{Y}^{i}$ contains the sink of the action on ${\mathcal{U}}_{Y}$. Then, given any point $x\in{\mathcal{U}}_{Y}$ there exists a connected chain of closures of orbits $C_{1}\cup\dots\cup C_{k}$ in ${\mathcal{U}}_{Y}$ such that $C_{1}$ meets the sink of ${\mathcal{U}}_{Y}$, the source of $C_{j}$ is the sink of $C_{j+1}$ for every $j$, and $x\in C_{k}$. By Lemma 4.9 there are no such chains joining two different components of ${\mathcal{U}}_{Y}^{i}$, hence ${\mathcal{U}}_{Y}^{i}$ is irreducible. To prove the last statement it is enough to notice that, for every $y\in Y$, every connected chain of closures of orbits meeting ${\mathcal{M}}_{y}$ is contained in ${\mathcal{M}}_{y}$. ∎
@@
If ${\mathcal{U}}_{Y}^{-}\not=\emptyset$ (resp. ${\mathcal{U}}_{Y}^{+}\not=\emptyset$) then it is irreducible, is the sink (resp. the source) of the action on ${\mathcal{U}}_{Y}$ and meets ${\mathcal{M}}_{y}$ for every $y\in Y$. Moreover, if ${\mathcal{U}}_{Y}^{0}$ is reducible, then ${\mathcal{U}}_{Y}^{\pm}$ are not empty, irreducible, and are the extremal fixed point components of the action.
@@
Consider the faithful action on ${\mathbb{P}}^{n}$ which has two linear spaces ${\mathbb{P}}^{k}$ ${\mathbb{P}}^{n-k-1}$ as extremal fixed point components $(1\leq k\leq n-2)$ and the induced action on the Grassmannian of lines ${\rm A}_{n}(2)$. This action has extremal fixed point components isomorphic to ${\rm A}_{k}(2)$ and ${\rm A}_{n-k-1}(2)$, and one inner fixed point $Y$ component isomorphic to ${\rm A}_{k}(1)\times{\rm A}_{n-k-1}(1)$; in this case ${\mathcal{U}}_{Y}^{0}$, which parametrizes the lines contained in $Y$, has two irreducible components.
@@
$p({\mathcal{U}}_{Y}^{-})=\bigsqcup_{\mu_{L}(Y_{i})=\mu_{L}(Y)-d}p({\mathcal{U}}_{Y_{i}}^{+})\cap p({\mathcal{U}}^{-}_{Y})$
@@ 5. Action on $X$: orbit graph and BB-cells
$X$ is a smooth projective variety of Picard number one, endowed with an equalized ${\mathbb{C}}^{*}$-action of criticality $r$, admitting a beautiful family of rational curves ${\mathcal{M}}$, with evaluation morphism $q:{\mathcal{U}}\to{\mathcal{M}}$ (see Section 2.7). We denote by $L$ the ample generator of $\mathop{\rm Pic}\nolimits(X)$, and by $d$ the $L$-degree of the curves of ${\mathcal{M}}$.
We start by describing the fixed point components of the action on ${\mathcal{U}}$, which, as noticed in Section 4.2, are contained in $\bigsqcup_{Y\in{\mathcal{Y}}}\,{\mathcal{U}}_{Y}$.
@@
Let $X$ and ${\mathcal{M}}$ be as in Setup 5.1. Let $Y\in{\mathcal{Y}}$ be a fixed point component of the ${\mathbb{C}}^{*}$-action on $X$, and set ${\mathcal{U}}_{Y}=q^{-1}(Y)$. Then:
@@
    If $Y$ is extremal and $\dim Y=0$, then the action on ${\mathcal{U}}_{Y}$ is trivial.
@@
    If $Y$ is the sink (resp. the source) and $\dim Y>0$, then ${\mathcal{U}}_{Y}^{0}$ and ${\mathcal{U}}_{Y}^{+}$ (resp. ${\mathcal{U}}_{Y}^{-}$ and ${\mathcal{U}}_{Y}^{0}$) are non-empty, and are the sink and the source of the action on ${\mathcal{U}}_{Y}$. There are no inner fixed point components.
@@
    If $Y$ is an inner fixed point component then ${\mathcal{U}}_{Y}^{-}$ and ${\mathcal{U}}_{Y}^{+}$ are non-empty, and are the sink and the source of the action on ${\mathcal{U}}_{Y}$. The inner fixed point components of the action are the irreducible components of ${\mathcal{U}}_{Y}^{0}$, if ${\mathcal{U}}_{Y}^{0}\not=\emptyset$.
@@
In case (e2), we have that ${\mathcal{U}}_{Y}^{-}=\emptyset$. We will show that ${\mathcal{U}}_{Y}^{0}\neq\emptyset$, from which Lemma 4.11, together with Lemma 4.9 will tell us that ${\mathcal{U}}_{Y}^{0}$, ${\mathcal{U}}_{Y}^{+}$ are, respectively, the sink and the source of the action. Since $\mathop{\rm Pic}\nolimits(X)\simeq{\mathbb{Z}}$, by [26, Lemma 2.8.1] the union of closures of the BB-cells $B^{-}(Y^{\prime})$ over the fixed point components $Y^{\prime}$ different from $Y$ has codimension at least two in $X$. In particular, by [17, Proposition II.3.7], the general curve $C$ of ${\mathcal{M}}$ is contained in its complement, which is the BB-cell $X^{-}(Y)$. It then follows that the limit curve $C_{-}$ must be contained in $Y$; hence ${\mathcal{U}}_{Y}^{0}\neq\emptyset$.
@@
    The weights of the action are $0<d<2d<\dots<(r-1)d<rd$.
@@
    For every $j=0,\dots,r$ there exists a unique $Y_{j}\in{\mathcal{Y}}$ such that $\mu_{L}(Y_{j})=jd$.
@@
    Through every point $y\in Y_{j}$ with $j\not=0$ there are bridges linking $Y_{j-1}$ and $Y_{j}$.
@@
Let $Y$ be a fixed point component; by Proposition 5.2, if $Y$ is the sink (resp. the source), then ${\mathcal{U}}_{Y}^{+}$ (resp. ${\mathcal{U}}_{Y}^{-}$) is non-empty, and if $Y$ is an inner fixed point component, then both ${\mathcal{U}}_{Y}^{-}$, ${\mathcal{U}}_{Y}^{+}$ are non-empty.
Moreover, by Corollary 4.12, if ${\mathcal{U}}_{Y}^{\pm}\not=\emptyset$, then ${\mathcal{U}}_{Y}^{\pm}\cap{\mathcal{M}}_{y}\not=\emptyset$ for every $y\in Y$. In particular, for every fixed point not contained in the sink there exists a bridge with source at that point and for every fixed point not contained in the source there exists a bridge with sink at that point. In particular, if $a_{i}$ is a weight with $0<i<r$ then also $a_{i}\pm d$ are weights of the action. Since every fixed point component can then be joined to the sink by a connected chain of bridges, by Lemma 2.3 every weight is a multiple of $d$.
It remains to prove the uniqueness in (2) and we proceed by contradiction. Let $j$ be the first index for which there exist two fixed point components $Y_{j},Y^{\prime}_{j}$ of weight $jd$. Since the sink and the source are the only fixed point components of weight 00 and $rd$, respectively, we have that $0<j<r$, so by the first part of the proof we have that the the unique component of weight $(j-1)d$ is linked by a bridge to both $Y_{j}$ and $Y^{\prime}_{j}$, but this contradicts Corollary 4.14. ∎

### v2

@@ 4. Induced actions on families of rational curves
$(X,L)$ is a smooth polarized pair, endowed with an equalized ${\mathbb{C}}^{*}$-action, and admitting a beautiful family of rational curves ${\mathcal{M}}$, of $L$-degree $d$, with evaluation morphism $q:{\mathcal{U}}\to{\mathcal{M}}$ (see Section 2.7).
@@ 4.1. Action on ${\mathcal{M}}$
Let $C$ be a curve such that $[C]\in{\mathcal{M}}$; the point $[C]$ is fixed by the induced action if and only if $C$ is a ${\mathbb{C}}^{*}$-invariant curve, that is, if either $C\subset Y$ with $Y\in{\mathcal{Y}}$, or $C$ is the closure of an orbit of the action. The latter curves will be called bridges of the family ${\mathcal{M}}$ (with respect to the ${\mathbb{C}}^{*}$-action). If a bridge $C$ has sink and source in fixed components $Y,Y^{\prime}$, we say that the bridge $C$ links $Y$ and $Y^{\prime}$. Note that in this case, by Lemma 2.3, $\mu_{L}(Y)\not=\mu_{L}(Y^{\prime})$, hence $Y\not=Y^{\prime}$.
@@
Given a non-fixed point $[C]\in{\mathcal{M}}$ we construct a rational ruled surface $S_{C}$ as follows: we take the closure of the orbit of $[C]$, $\overline{{\mathbb{C}}^{*}[C]}$, which is a rational curve, we take its normalization $\nu:{\mathbb{P}}^{1}\to\overline{{\mathbb{C}}^{*}[C]}$ and we consider the fiber product:
@@
By construction, the surface $S_{C}$ is isomorphic to $\mathbb{F}_{e}:={\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{1}}\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(e))$ for some $e\geq 0$, and is endowed with a ${\mathbb{C}}^{*}$-action with two invariant fibers $F_{-}$, $F_{+}$, which are mapped via $\overline{q}$ to the limit curves of $C$.
@@
Note that $\overline{q}(S_{C})$ has dimension two, so $\overline{q}$ can contract at most one curve in $S_{C}$, otherwise we would have a one-dimensional family of curves of ${\mathcal{M}}$ passing by two points of $X$; this, by Bend-and-Break (see [7, Proposition 3.2]) would contradict the fact that ${\mathcal{M}}$ is an unsplit family. If the curve $C$ contains a fixed point of the action, then $S_{C}$ contains a section which is contracted by $\overline{q}$.
@@
We will now use Construction 1 to prove that the equalization of the action on $X$ implies the equalization of the induced ${\mathbb{C}}^{*}$-action on ${\mathcal{M}}$.
@@
If $e=0$, we consider the ${\mathbb{C}}^{*}$-action on the space of global sections $\mathop{\rm H}\nolimits^{0}({\mathbb{P}}^{1},{\mathcal{E}})$, that has two linearly independent eigenvectors, providing two ${\mathbb{C}}^{*}$-invariant surjective morphisms ${\mathcal{E}}\to{\mathcal{O}}_{{\mathbb{P}}^{1}}$, that correspond to two ${\mathbb{C}}^{*}$-invariant sections of $S_{C}\to{\mathbb{P}}^{1}$, denoted by $C_{0},C_{1}$. We conclude by noting that, by Remark 4.3, $\overline{q}$ cannot contract both $C_{0}$ and $C_{1}$.
@@
We finish the section by describing the possibilities for the surface $S_{C}$ and the limit curves $C_{-},C_{+}$.
@@
    $e=1$, one among $C_{-}$ and $C_{+}$ is fixed, and the other is a bridge;
@@
Set ${\mathcal{E}}:={\mathcal{O}}_{{\mathbb{P}}^{1}}(1)\oplus{\mathcal{O}}_{{\mathbb{P}}^{1}}(e+1)$ and let $a_{-},b_{-}$ (resp. $a_{+},b_{+}$) be the weights of the action at ${\mathcal{E}}_{[C_{-}]}$ (resp. at ${\mathcal{E}}_{[C_{+}]}$). By [26, Lemma 2.16], up to renaming and normalizing them, we have $a_{-}=0,a_{+}=1$ and $b_{+}=b_{-}+e+1$. Moreover, since the action on $F_{-}$ and $F_{+}$ is equalized, we have that $|b_{-}|\leq 1$ and $0\leq b_{+}\leq 2$. We also note that the fiber $F_{-}$ (resp. the fiber $F_{+}$) is a fixed point component if and only if $b_{-}=0$ (resp. $b_{+}=1$). Therefore we have the following possibilities, which lead to the cases listed in the statement:
@@ 00 $2$ $1$ fixed orbit
If $C$ contains a fixed point $y$, then every fiber of $S_{C}$ over ${\mathbb{P}}^{1}$ contains a point mapping to $y$. In other words, $S_{C}$ contains a section over ${\mathbb{P}}^{1}$ contracted via $\overline{q}$ to $y$. Since $C$ is not fixed, we may conclude that $e>0$. ∎
@@
Figure 1. The possible cases in Proposition 4.6.
@@
  • in case 1), either $C_{+}$ links $Y$ to a component $Y^{\prime}$ of $L$-weight $\mu_{L}(Y)+d$, or $C_{-}$ links $Y$ to a component $Y^{\prime}$ of $L$-weight $\mu_{L}(Y)-d$;
@@ 4.2. Action on ${\mathcal{U}}$
Let $u\in{\mathcal{U}}$ be a non-fixed point, and let $G_{u},G_{p(u)},G_{q(u)}\subset{\mathbb{C}}^{*}$ be the isotropy subgroups of $u\in{\mathcal{U}}$, $p(u)\in{\mathcal{M}}$, $q(u)\in X$. Clearly we have that $G_{u}\subset G_{p(u)}\cap G_{q(u)}$. Since the closure of ${\mathbb{C}}^{*}u$ cannot be contracted by both $p$ and $q$, we have that either $G_{p(u)}$ or $G_{q(u)}$ are finite. Moreover, since the actions of ${\mathbb{C}}^{*}$ on $X$ and ${\mathcal{M}}$ are equalized, we have that either $G_{p(u)}$ or $G_{q(u)}$ are trivial, and we conclude that $G_{u}$ is trivial as well. ∎
@@
Since $p$ is equivariant, the image of a fixed point in ${\mathcal{U}}_{Y}$ must be a point $[C]$ such that either $C\subset Y$, or $C$ is a bridge. In other words, the fixed point locus of ${\mathcal{U}}_{Y}$ can be written as a disjoint union of (a priori reducible) closed subsets
@@
Note that ${\mathcal{U}}_{Y}=\bigsqcup_{y\in Y}{\mathcal{U}}_{y}$, and that every ${\mathcal{U}}_{y}$ is ${\mathbb{C}}^{*}$-invariant. In particular, every orbit in ${\mathcal{U}}_{Y}$ is contained in ${\mathcal{U}}_{y}$ for some $y\in Y$. The following Lemma, which is a straightforward consequence of Corollary 4.7, describes the ${\mathbb{C}}^{*}$-orbits on ${\mathcal{U}}_{Y}$:
@@
If  ${\mathcal{U}}_{Y}^{i}$ (for $i=-,0,+$) contains an extremal fixed point component, then it is irreducible, and meets ${\mathcal{U}}_{y}$ for every $y$.
@@
Assume, without loss of generality, that ${\mathcal{U}}_{Y}^{i}$ contains the sink of the action on ${\mathcal{U}}_{Y}$. Then, given any point $x\in{\mathcal{U}}_{Y}$ there exists a connected chain of closures of orbits $C_{1}\cup\dots\cup C_{k}$ in ${\mathcal{U}}_{Y}$ such that $C_{1}$ meets the sink of ${\mathcal{U}}_{Y}$, the source of $C_{j}$ is the sink of $C_{j+1}$ for every $j$, and $x\in C_{k}$. By Lemma 4.9 there are no such chains joining two different components of ${\mathcal{U}}_{Y}^{i}$, hence ${\mathcal{U}}_{Y}^{i}$ is irreducible. To prove the last statement it is enough to notice that, for every $y\in Y$, every connected chain of closures of orbits meeting ${\mathcal{U}}_{y}$ is contained in ${\mathcal{U}}_{y}$. ∎
@@
If ${\mathcal{U}}_{Y}^{-}\not=\emptyset$ (resp. ${\mathcal{U}}_{Y}^{+}\not=\emptyset$), then it is irreducible, it is the sink (resp. the source) of the action on ${\mathcal{U}}_{Y}$, and meets ${\mathcal{U}}_{y}$ for every $y\in Y$. Moreover, if ${\mathcal{U}}_{Y}^{0}$ is reducible, then ${\mathcal{U}}_{Y}^{\pm}$ are not empty, irreducible, and are the extremal fixed point components of the action.
@@
Consider the equalized action on ${\mathbb{P}}^{n}$ which has two linear spaces ${\mathbb{P}}^{k}$ ${\mathbb{P}}^{n-k-1}$ as extremal fixed point components $(1\leq k\leq n-2)$ and the induced action on the Grassmannian of lines ${\rm A}_{n}(2)$. This action has extremal fixed point components isomorphic to ${\rm A}_{k}(2)$ and ${\rm A}_{n-k-1}(2)$, and one inner fixed point $Y$ component isomorphic to ${\rm A}_{k}(1)\times{\rm A}_{n-k-1}(1)$; in this case ${\mathcal{U}}_{Y}^{0}$, which parametrizes the lines contained in $Y$, has two irreducible components.
@@
$p({\mathcal{U}}_{Y}^{-})=\bigsqcup_{\mu_{L}(Y_{i})=\mu_{L}(Y)-d}p({\mathcal{U}}_{Y_{i}}^{+})\cap p({\mathcal{U}}^{-}_{Y}),$
@@ 5. Action on $X$: orbit graph and BB-cells
$X$ is a smooth projective variety of Picard number one, endowed with an equalized ${\mathbb{C}}^{*}$-action of criticality $r$, admitting a beautiful family of rational curves ${\mathcal{M}}$, with evaluation morphism $q:{\mathcal{U}}\to X$ (see Section 2.7). We denote by $L$ the ample generator of $\mathop{\rm Pic}\nolimits(X)$, and by $d$ the $L$-degree of the curves of ${\mathcal{M}}$.
We start by describing the fixed point components of the action on ${\mathcal{U}}$, which, as noticed in Section 4.2, are contained in $\bigsqcup_{Y\in{\mathcal{Y}}}\,{\mathcal{U}}_{Y}$, where ${\mathcal{U}}_{Y}=q^{-1}(Y)$. For every $Y\in{\mathcal{Y}}$ we will consider the fixed point components ${\mathcal{U}}_{Y}^{-},{\mathcal{U}}_{Y}^{0},{\mathcal{U}}_{Y}^{+}$, as defined in Section 4.2.
@@
Let $X$ and ${\mathcal{M}}$ be as in Setup 5.1. Let $Y\in{\mathcal{Y}}$ be a fixed point component of the ${\mathbb{C}}^{*}$-action on $X$. Then:
@@
    if $Y$ is extremal and $\dim Y=0$, then the action on ${\mathcal{U}}_{Y}$ is trivial;
@@
    if $Y$ is the sink (resp. the source) and $\dim Y>0$, then ${\mathcal{U}}_{Y}^{0}$ and ${\mathcal{U}}_{Y}^{+}$ (resp. ${\mathcal{U}}_{Y}^{-}$ and ${\mathcal{U}}_{Y}^{0}$) are non-empty, and are the sink and the source of the action on ${\mathcal{U}}_{Y}$. There are no inner fixed point components;
@@
    if $Y$ is an inner fixed point component then ${\mathcal{U}}_{Y}^{-}$ and ${\mathcal{U}}_{Y}^{+}$ are non-empty, and are the sink and the source of the action on ${\mathcal{U}}_{Y}$; the inner fixed point components of the action are the irreducible components of ${\mathcal{U}}_{Y}^{0}$, if ${\mathcal{U}}_{Y}^{0}\not=\emptyset$.
@@
In case (e2), we have that ${\mathcal{U}}_{Y}^{-}=\emptyset$. We will show that ${\mathcal{U}}_{Y}^{0}\neq\emptyset$; then, by Lemma 4.11 and Lemma 4.9 we will get that ${\mathcal{U}}_{Y}^{0}$, ${\mathcal{U}}_{Y}^{+}$ are, respectively, the sink and the source of the action. Since $\mathop{\rm Pic}\nolimits(X)\simeq{\mathbb{Z}}$, by [26, Lemma 2.8.1] the union of closures of the BB-cells $B^{-}(Y^{\prime})$ over the fixed point components $Y^{\prime}$ different from $Y$ has codimension at least two in $X$. In particular, by [17, Proposition II.3.7], the general curve $C$ of ${\mathcal{M}}$ is contained in its complement, which is the BB-cell $X^{-}(Y)$. It then follows that the limit curve $C_{-}$ must be contained in $Y$; we conclude that ${\mathcal{U}}_{Y}^{0}\neq\emptyset$.
@@
    the weights of the action are $0<d<2d<\dots<(r-1)d<rd$;
@@
    for every $j=0,\dots,r$ there exists a unique $Y_{j}\in{\mathcal{Y}}$ such that $\mu_{L}(Y_{j})=jd$;
@@
    through every point $y\in Y_{j}$ with $j\not=0$ there are bridges linking $Y_{j-1}$ and $Y_{j}$;
@@
Let $Y$ be a fixed point component. By Proposition 5.2, if $Y$ is the sink (resp. the source), then ${\mathcal{U}}_{Y}^{+}$ (resp. ${\mathcal{U}}_{Y}^{-}$) is non-empty, and if $Y$ is an inner fixed point component, then both ${\mathcal{U}}_{Y}^{-}$, ${\mathcal{U}}_{Y}^{+}$ are non-empty.
Moreover, by Corollary 4.12, if ${\mathcal{U}}_{Y}^{\pm}\not=\emptyset$, then ${\mathcal{U}}_{Y}^{\pm}\cap{\mathcal{U}}_{y}\not=\emptyset$ for every $y\in Y$. In particular, for every fixed point not contained in the sink there exists a bridge with source at that point, and for every fixed point not contained in the source there exists a bridge with sink at that point. Therefore, if $a_{i}$ is a weight with $0<i<r$ then also $a_{i}\pm d$ are weights of the action. Since every fixed point component can be joined to the sink by a connected chain of bridges, by Lemma 2.3 every weight is a multiple of $d$.
It remains to prove the uniqueness in (2) and we proceed by contradiction. Let $j$ be the minimum index for which there exist two fixed point components $Y_{j},Y^{\prime}_{j}$ of weight $jd$. Since the sink and the source are the only fixed point components of weight 00 and $rd$, respectively, we have that $0<j<r$. Hence, by the first part of the proof, we have that the the unique component of weight $(j-1)d$ is linked by a bridge to both $Y_{j}$ and $Y^{\prime}_{j}$; this contradicts Corollary 4.14. ∎